# Crossings and diffusion in Poisson driven marked random connection models

### v1

@@ 6.3 Probabilistic bounds
The next lemma is an application of Lemmas 6.2 and 6.3 following [Chayes86]*Lemma 3.3.
@@
  • going from $s=0$ to $s=M-1$ the probability that $x^{s}_{1}$ is occupied, conditioned to the revealed status (i.e. occupied or not occupied) of $x_{1}^{s^{\prime}}$ with $s^{\prime}<s$, is at least $p$,
  • every time we check which one of the two cases in (36) is valid, the probability that $x^{s}_{j+1}$ is linked to $x_{k}^{s}$, conditioned to the accumulated information (i.e. to the knowledge of $C^{a}_{1}$ with $10\leq a\leq M-$ 1, of $C_{r}^{s^{\prime}}$ with $0\leq s^{\prime}<s$ and $1<r\leq\sharp\Lambda_{L}^{\prime}$ and of $C_{r}^{s}$ with $1<r\leq j$), is at least $p$.
@@
By standard arguments (cf. e.g. [Chayes86]*Lemma 3.3, [Faggionato21]*Proposition 3.6), in order to prove Theorem 2.4, it is enough to show the following: for any $\lambda^{\prime}>\lambda$, there exist $c_{1},c_{2}>0$ such that, for all ${\ell}\geq 1$, we have
$P_{\lambda^{\prime},\nu}\left(\mathcal{M}_{\ell}\geq c_{1}\ell\right)\geq 1-\exp\left(-c_{2}\,\ell\right)\,,$ (37)
@@
$\Lambda_{6N}(-12Ne_{1})$$\Lambda_{6N}$$\Lambda^{0}$$\Lambda^{1}$$\Lambda^{2}$$\Lambda^{3}$$\Lambda$$S_{0}$$\mathcal{S}^{1}$$\mathcal{S}^{3}$$\mathcal{S}^{2}$
Figure 1: Illustration of Definition 7.1. The last $d-2$ dimensions and marks are suppressed.
@@
Choosing the scales in (23) suitably, it holds that $P(\text{$(0,0)$ is occupied})>p_{0}$.
@@
Fix $u>0$ to be chosen later. We apply Lemma 5.1 with $R:=\Lambda_{N}(z_{1})\times{\mathbb{M}}$, $A:=\bar{S}$, $B:=(z_{1}+F^{1,\sigma^{1}}_{N})\times{\mathbb{M}}$ with $\sigma^{1}$ the sequence of the signs of the coordinates of $-z_{1}$ (exceptionally, in this section we define the sign of 00 as $+1$), $\tilde{\mathbb{X}}:=\xi_{1}\cap R$ under $P_{1}$ (and therefore under $P$), ${\mathbb{Y}}:=(\xi\setminus\bar{S})\cap R$ under $P_{1}$, ${\mathbb{X}}:=\xi\cap R$ under $P$. By Lemmas 5.1 and 5.5 we then get
@@
We repeat the same reasoning 4 more times (using a new $\xi_{i}$ each time) until we obtain a seed $\mathcal{S}_{6}\subset\Lambda_{n-K}(z_{6})\times{\mathbb{M}}_{g}\subset[-N/2-6\sqrt{N},6\sqrt{N}]\times[-4N-6\sqrt{N},4N+6\sqrt{N}]^{d-1}\times{\mathbb{M}}_{g}$. Note that the projection on the first coordinate of $\Lambda_{n-K}(z_{6})$ is included in $[-2N,2N]$ (by taking $N$ large). We then seek to adjust the other coordinates by repeating the same procedure another $4(d-1)$ times. We start with the second coordinate. For $i\in\{6,7,8,9\}$ we try to connect $S_{i}$ with $z_{i}+F_{N}^{2,\sigma^{i}}$ where $\sigma^{i}$ is the sign sequence of $-z_{i}$, thus leading to the seed $\mathcal{S}_{i+1}\subset\Lambda_{n-K}(z_{i+1})\times{\mathbb{M}}_{g}$. Then we do the same for the third coordinate and so on. Notice that, at each step when adjusting the $j$–th coordinate, $|z_{i+1}(j)|\leq\max(|z_{i}(j)|-3N/4,3N/2)$, while $|z_{i+1}(j^{\prime})|\leq\max(|z_{i}(j^{\prime})|+\sqrt{N},3N/2)$ for $j^{\prime}\in\{1,\dots,d\}\setminus\{j\}$.
@@
From this point, for each $i\in\{1,2,3\}$ we proceed similarly going from $\mathcal{S}_{4d+2}$ towards $\Lambda^{i}\times{\mathbb{M}}$. For concreteness, consider $i=2$, the others being analogous. Further for concreteness, assume that $z_{4d+2}\in[N-\sqrt{N},2N]\times[-2N-3\sqrt{N},2N+3\sqrt{N}]^{d-1}$ (if this is not the case, up to three additional steps are performed to reach this case, setting $\sigma_{1}=1$ rather than the sign of the first coordinate of $-z_{j}$). We perform another 3 steps to the right, obtaining $z_{4d+5}\in[4N-4\sqrt{N},5N+3\sqrt{N}]\times[-2N-6\sqrt{N},2N+6\sqrt{N}]^{d-1}$. We now consider three cases for the first coordinate $\zeta$ of $z_{4d+5}$.
If $\zeta\in[9N/2+2\sqrt{N},5N-2\sqrt{N}]$, then an additional step as before yields the desired seed $\mathcal{S}_{4d+6}\subset\Lambda^{2}$ and we are done. Assume that $\zeta\in[4N-4\sqrt{N},9N/2+2\sqrt{N}]$ and perform a step as before to get that the first coordinate of $z_{4d+6}$ belongs to $[5N-5\sqrt{N},11N/2+3\sqrt{N}]$. At this point, we need more care, since $\Lambda_{N}(z_{4d+6})\not\subset\Lambda$. However, we may proceed in the same way as above, setting $R=\Lambda_{N}(z_{4d+6})\cap((-\infty,23N/4]\times{\mathbb{R}}^{d-1})\subset\Lambda$ and replacing $z_{4d+6}+F^{1,\sigma^{4d+6}}_{N}$ by $F^{\prime}=R\cap([23N/4-1,23N/4]\times{\mathbb{R}}^{d-1})$. Indeed, for any seed $S\subset\Lambda_{n-K}$ and $x\in[n+1,N]$, we have
@@
In total, in at most $264d+$ 26 steps, we have obtained that
@@ Acknowledgements
The authors thank Lorenzo Dello Schiavo and Markus Heydenreich for helpful discussions.
@@ Appendix A The effective homogenized matrix $D(\rho)$
To each $\zeta\in\mathcal{N}({\mathbb{W}})$ we associate a family of edges in ${\mathbb{R}}^{d}$ denoted by $E(\zeta)$, given by the pairs $\{x,y\}$ such that for some ($(\{t,t^{\prime}\},u)$ it holds $x=\pi(t)$, $y=\pi(t^{\prime})$ and $u\leq\varphi(t,t^{\prime})$. If the graph with the above edge set $E(\zeta)$ (and vertices their endpoints) has a unique infinite cluster, we denote by $\mathcal{C}(\zeta)$ this cluster, otherwise we set $\mathcal{C}(\zeta):=\varnothing$. Finally, we define $\mathcal{E}(\zeta)$ as the family of edges $\{x,y\}\in E(\zeta)$ with vertices $x,y$ in $\mathcal{C}(\zeta)$.
We introduce the simple point process (SPP) $\mathcal{N}({\mathbb{W}})\ni\zeta\to\hat{\zeta}\in\mathcal{N}({\mathbb{R}}^{d})$ defined as $\hat{\zeta}:=\mathcal{C}(\zeta)$ and the conductance field $c(x,y,\zeta)$ defined as $c(x,y,\zeta):=\mathbbm{1}(\{x,y\}\in E(\zeta))$. Then the weighted graph in [Faggionato25] associated to the above SPP and conductance field coincides in law to the infinite cluster of the RCM$(\rho,\nu,\varphi)$.
The Abelian group ${\mathbb{G}}={\mathbb{R}}^{d}$ acts on the Euclidean space ${\mathbb{R}}^{d}$ by the translations $(\tau_{g})_{g\in{\mathbb{G}}}$, where $\tau_{g}x:=x+g$ for $x\in{\mathbb{R}}^{d}$. ${\mathbb{G}}$ acts on on $\mathcal{N}({\mathbb{W}})$ by the action $(\theta_{g})_{g\in{\mathbb{G}}}$ where, given $\zeta\in\mathcal{N}({\mathbb{W}})$, the set $\theta_{g}\zeta$ is given by the elements ($\left(\{(x-g,m),(x^{\prime}-g,m^{\prime})\},u\right)$ as ($\left(\{(x,m),(x^{\prime},m^{\prime})\},u\right)$ varies in $\zeta$. Note that $\mathcal{P}_{\rho}$ is stationary and ergodic for the action $(\theta_{g})_{g\in{\mathbb{G}}}$. Indeed $\mathcal{P}_{\rho}$, is even mixing (the proof of mixing is simplified by Assumption 2.3(ii)).

### v2

@@ 6.3 Probabilistic bounds
The next lemma is an application of Lemmas 6.2 and 6.3 following [Chayes86]*Lemma $3.$
@@
  • going from $s=0$ to s=M1s=M-1 the probability that $x^{s}_{1}$ is occupied, conditioned to the revealed status (i.e. occupied or not occupied) of $x_{1}^{s^{\prime}}$ with $s^{\prime}<s$, is at least $p$,
  • every time we check which one of the two cases in (36) is valid, the probability that $x^{s}_{j+1}$ is linked to $x_{k}^{s}$, conditioned to the accumulated information (i.e. to the knowledge of $C^{a}_{1}$ with $0\leq a\leq M-1$, of $C_{r}^{s^{\prime}}$ with $0\leq s^{\prime}<s$ and $1<r\leq\sharp\Lambda_{L}^{\prime}$ and of $C_{r}^{s}$ with $1<r\leq j$), is at least $p$.
@@
By standard arguments (cf. e.g. [Chayes86]*Lemma 3.3, [Faggionato21]*Proposition 3.6), in order to prove Theorem 2.4, it is enough to show the following: for any $\lambda^{\prime}>\lambda$, there exist $c_{1},c_{2}>0$ such that, for all $\ell$ large enough, we have
$P_{\lambda^{\prime},\nu}\otimes Q\left(\mathcal{M}_{\ell}\geq c_{1}\ell\right)\geq 1-\exp\left(-c_{2}\,\ell\right)\,$, (37)
@@
[…]
Figure 1: Illustration of Definition 7.1. The last d2d-2 dimensions and marks are suppressed.
@@
[…]
@@
Fix $u>0$ to be chosen later. We apply Lemma 5.1 with $R:=\Lambda_{N}(z_{1})\times{\mathbb{M}}$, $A:=\bar{S}$, $B:=(z_{1}+F^{1,\sigma^{1}}_{N})\times{\mathbb{M}}$ with $\sigma^{1}$ the sequence of the signs of the coordinates of $z_{1}$ (exceptionally, in this section we define the sign of 0 as +1+1), $\tilde{\mathbb{X}}:=\xi_{1}\cap R$ under $P_{1}$ (and therefore under $P$), ${\mathbb{Y}}:=(\xi\setminus\bar{S})\cap R$ under $P_{1}$, ${\mathbb{X}}:=\xi\cap R$ under $P$. By Lemmas 5.1 and 5.5 we then get
@@
We repeat the same reasoning 4 more times (using a new $\xi_{i}$ each time) until we obtain a seed $\mathcal{S}_{6}\subset\Lambda_{n-K}(z_{6})\times{\mathbb{M}}_{g}\subset[-N/2-6\sqrt{N},6\sqrt{N}]\times[-4N-6\sqrt{N},4N+6\sqrt{N}]^{d-1}\times{\mathbb{M}}_{g}$. Note that the projection on the first coordinate of $\Lambda_{n-K}(z_{6})$ is included in [2N,2N][-2N,2N] (by taking $N$ large). We then seek to adjust the other coordinates by repeating the same procedure another 4(d1)4(d-1) times. We start with the second coordinate. For $i\in\{6,7,8,9\}$ we try to connect $S_{i}$ with $z_{i}+F_{N}^{2,\sigma^{i}}$ where $\sigma^{i}$ is the sign sequence of $z_{i}$, thus leading to the seed $\mathcal{S}_{i+1}\subset\Lambda_{n-K}(z_{i+1})\times{\mathbb{M}}_{g}$. Then we do the same for the third coordinate and so on. Notice that, at each step when adjusting the jj–th coordinate, $|z_{i+1}(j)|\leq\max(|z_{i}(j)|-3N/4,3N/2)$, while $|z_{i+1}(j^{\prime})|\leq\max(|z_{i}(j^{\prime})|+\sqrt{N},3N/2)$ for $j^{\prime}\in\{1,\dots,d\}\setminus\{j\}$.
@@
From this point, for each $i\in\{1,2,3\}$ we proceed similarly going from $\mathcal{S}_{4d+2}$ towards $\Lambda^{i}\times{\mathbb{M}}$. For concreteness, consider $i=2$, the others being analogous. Further for concreteness, assume that $z_{4d+2}\in[N-\sqrt{N},2N]\times[-2N-3\sqrt{N},2N+3\sqrt{N}]^{d-1}$ (if this is not the case, up to three additional steps are performed to reach this case, setting $\sigma_{1}=1$ rather than the sign of the first coordinate of $z_{j}$). We perform another 3 steps to the right, obtaining $z_{4d+5}\in[4N-4\sqrt{N},5N+3\sqrt{N}]\times[-2N-6\sqrt{N},2N+6\sqrt{N}]^{d-1}$. We now consider three cases for the first coordinate $\zeta$ of $z_{4d+5}$.
If $\zeta\in[9N/2+2\sqrt{N},5N-2\sqrt{N}]$, then an additional step as before yields the desired seed $\mathcal{S}_{4d+6}\subset\Lambda^{2}$ and we are done. Assume that $\zeta\in[4N-4\sqrt{N},9N/2+2\sqrt{N}]$ and perform a step as before to get that the first coordinate of $z_{4d+6}$ belongs to $5N-5\sqrt{N},11N/2+3\sqrt{N}]$. At this point, we need more care, since $\Lambda_{N}(z_{4d+6})\not\subset\Lambda$. However, we may proceed in the same way as above, setting $R=\Lambda_{N}(z_{4d+6})\cap((-\infty,23N/4]\times{\mathbb{R}}^{d-1})\subset\Lambda$ and replacing $z_{4d+6}+F^{1,\sigma^{4d+6}}_{N}$ by $F^{\prime}=R\cap([23N/4-1,23N/4]\times{\mathbb{R}}^{d-1})$. Indeed, for any seed $S\subset\Lambda_{n-K}$ and $x\in[n+1,N]$, we have
@@
In total, in at most $4d+26$ steps, we have obtained that
@@ Acknowledgements
The authors thank Lorenzo Dello Schiavo, Markus Heydenreich and Mathew Penrose for helpful discussions.
@@ Appendix A The effective homogenized matrix $D(\rho)$
To each $\zeta\in\mathcal{N}({\mathbb{W}})$ we associate a family of edges in ${\mathbb{R}}^{d}$ denoted by $E(\zeta)$, given by the pairs $\{x,y\}$ such that for some $(\{t,t^{\prime}\},u)$ it holds $x=\pi(t)$, $y=\pi(t^{\prime})$ and $u\leq\varphi(t,t^{\prime})$. If the graph with the above edge set $E(\zeta)$ (and vertices their endpoints) has a unique infinite cluster, we denote by $\mathcal{C}(\zeta)$ this cluster, otherwise we set $\mathcal{C}(\zeta):=\varnothing$. Finally, we define $\mathcal{E}(\zeta)$ as the family of edges $\{x,y\}\in E(\zeta)$ with vertices $x,y$ in $\mathcal{C}(\zeta)$.
[…]
The Abelian group ${\mathbb{G}}={\mathbb{R}}^{d}$ acts on the Euclidean space ${\mathbb{R}}^{d}$ by the translations $(\tau_{g})_{g\in{\mathbb{G}}}$, where $\tau_{g}x:=x+g$ for $x\in{\mathbb{R}}^{d}$. $\mathbb{G}$} acts on on $\mathcal{N}({\mathbb{W}})$ by the action $(\theta_{g})_{g\in{\mathbb{G}}}$ where, given $\zeta\in\mathcal{N}({\mathbb{W}})$, the set $\theta_{g}\zeta$ is given by the elements $\left(\{(x-g,m),(x^{\prime}-g,m^{\prime})\},u\right)$ as $\left(\{(x,m),(x^{\prime},m^{\prime})\},u\right)$ varies in $\zeta$. Note that $\mathcal{P}_{\rho}$ is stationary and ergodic for the action $(\theta_{g})_{g\in{\mathbb{G}}}$. Indeed $\mathcal{P}_{\rho}$, is even mixing (the proof of mixing is simplified by Assumption 2.3(ii)).